OpenStudy (anonymous):

How do I solve this? (51+17^2)/10

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

square the 17 add the 51 then divide by 10

OpenStudy (anonymous):

I know that, but there is an easier way? How would I deal with large complex numbers if I can't use a calculator?

OpenStudy (anonymous):

@Mertsj Any suggestions?

OpenStudy (mertsj):

You could always use logs.

OpenStudy (mertsj):

Or: \[\frac{51}{10}+\frac{17^2}{10}\]

OpenStudy (raden):

this is an easy way : (51+17^2)/10 = 17(3+17)/10 = 17(20)/10 = 34